OpenStudy (anonymous):

Factor completely 2x^3+6x^2+4x+12

OpenStudy (anonymous):

2x^3 + 4x^2 + 6x + 12 = 2 (x^3 + 2x^2 + 3x + 6) = 2 ( ( x^3 + 3x ) + (2x^2 + 6 ) ) = 2 ( x ( x^2 + 3 ) + 2 (x^2 + 3 ) ) = 2 ( x + 2 ) ( x^2 + 3 )